Historical data show that customers who download music from a popular Web service spend approxi-mately $26 per month, with a standard deviation of $8, normally distributed, but never spend more than $30. Assuming that each customer is independent of the others, predict the average amount of the next ten transactions using excel. (Hint: Develop a method to ensure that the random variate is limited to $30.)
Answer #1

SOLUTION :

from given data :

Mean = 26

Standard deviation = 8

P(X > 30) = P((x-mu)/s > (30 - 26)/8)

= P(z > 4/8)

= P(z > 0.5)

= 0.3085375 [since from z table]

= 0.3085

P(X > 30) = 0.3085.